OpenStudy (anonymous):

solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + y = 4 x + 3y = 22

OpenStudy (anonymous):

\[-3x+y=4\\x+3y=22\]First, we're going to rearrange the first to isolate a variable.\[y=3x+4\]Next, we substitute for the variable to eliminate it.\[x+3(3x+4)=22\\x+9x+12=22\\10x=10\\x=1\]Now, to find out the variable we eliminated.\[y=3(1)+4=7\]Therefore we know the two lines intersect at:\[(1,7)\]

OpenStudy (anonymous):

ok